# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

### v1

@@ Abstract
Ebola virus disease is a severe, often deadly form of hemorrhagic fever, and is rapidly transmissible. It’s spread through human contact is increased through the direct exchange of infected blood, body fluids, and organs, as well as contact with surfaces on which these infected fluids have been dropped. Ebola’s complicated dynamics spread beyond integer order models, representing a central intersection for the application of machine learning to control theory. In the present study, we construct an innovative prime framework of fractional order calculus which integrates data-driven machine learning with optimal control in order to model and analyze the Ebola virus disease. We develop an eight-compartment model based on the Caputo fractional derivative to incorporate the non-Markovian memory effects inherent in the disease’s progression and regression. The system explicitly structures the distinct transmission pathways from symptomatic and asymptomatic infections, as well as the critical transmission dynamics associated with deceased individuals. We analytically prove the global well-posedness of the model, compute the basic reproduction number $\mathcal{R}_{0}$ to characterize the epidemic threshold, and prove global stability theorems. Sensitivity analysis indicates that the basic reproduction number is most critically sensitive to transmission rate, incubation period, and the relative infectivity of the deceased. The treatment-safe burial synergy optimization control provides a simultaneous 86.5% of the target morbidity and mortality rate control of the population. The safe burial host intervention is the most effective control. Our study categorizes the best control strategies and orders them in terms of effectiveness. The proposed disease-informed neural network achieves practically perfect predictive accuracy and describe the test set $R^{2}$ values of 0.991 - 0.999, and correspondingly 99.1% - 99.9% prediction accuracy as well as align with the real-world epidemic behavior. The first successful combination of process-based modeling with artificial intelligence for dealing an outbreak of an infectious disease has been established in this work. This work presents an innovative mechanistic modeling framework for infectious diseases, which has the potential to advance epidemic response analysis.
Keywords: Fractional Calculus, Ebola Modeling, Epidemic Threshold ($\mathcal{R}_{0}$), Sensitivity Analysis, Optimal Control, Disease-Informed Neural Networks.
@@ 1 Introduction
Ebola Virus Disease (EVD) continues to be one of the most difficult problems for the world health organization to tackle, with estimated case fatality rates varying from concerning 25% to 90% across various outbreaks [1]. The Ebola Virus Disease (EVD) is classified as a tropical zoonosis which involves cross transmission from animal reservoirs to human populations [2, 3]. The earliest recorded case of the disease occurred in Zaire (now known as Democratic Republic of Congo) in 1976, named after the Ebola River where the first outbreak took place [4]. With data around confirmed cases and deaths standing at 28,646 and 11,323, respectively, and spanning through three countries; Liberia, Guinea, and Sierra Leone, the ‘West African Ebola Virus Disease epidemic’ (2014-16) is considered the biggest ‘Ebola’ outbreak in history. The surprising number of cases and deaths during the outbreak and the new ways of fighting the complexity both demonstrated the failure of prediction and outbreak control traditional prediction and control methods used [5]. The outbreak still emphasized the complexity controlled in the actual mathematics and logic of the base prediction and control frameworks prediction, controlling, and modeling control frameworks outlined around the multi and complex dimensions in the spread of the EVD. Mathematical techniques in epidemiology and their application span new prospects and new dimensions in the field [6, 7]. While these techniques were developed in pure science, their fundamental calculations and distributions resource divisions and practical works approximations [8]. The epidemiology of the western African case outbreak hence has triggered the much-needed efficiency flow in prediction case numbers and evaluating the efficiency of preventive measures and their real-time value [9]. The greatest puzzling challenge suffering in the environmental and acoustic scales spread while modalities of EVD transmissions is the combination of factors that various phoneme extents growing paradigm [10]. The epidemiology of Ebola Virus Disease (EVD) has many distinguishing features which imply that estimating the disease outbreaks remaining traditional techniques will not be sufficient. For instance, the disease’s incubation period is highly variable and spans anywhere from 2 to 21 days, with the 8-10 days interval more commonly renowned [11]. There is a further complication within the disease’s clinical presentation and development as initial fever, weakness, and diarrhea unaccompanied by bleeding and other clinical symptoms do not signal an active infection and contagiousness. Additionally, contact with infected animals and the consumption of bush-meat are the first chance for the disease’s human transmission from its natural reservoirs of fruit bats [12, 13]. Subsequently, the disease transmission post- infection is unlinked and random. Furthermore, symptomatic ($I_{s}$), asymptomatic ($I_{a}$), and post-mortem ($D$) transmission routes for the human population. [14]. The high level of infectiousness of dead body, coupled with the perseverance of customarily traditional burial practices in the region, requires mathematically sophisticated treatment beyond typical integer-order differential equations, as outlined in earlier works [15]. Evidence shows that washing the deceased as part of ethnically prescribed performs is a relevant vector for transmission. Besides, the prospect of reinfection with EVD remains very real, as survivors own only quickly protective, low-potency antibodies, with research showing the complete absence of Ebola antibodies in some survivor’s years post-recovery [16, 17]. Such an immunological situation is certainly an additional complication from the modeling perspective whereby the absence of an unified framework for archiving immune waning over a long period rightly remains a gap. The natural, spatial spread of EVD transmission improves so far, another level of complexity to the analytical modelling of the disease dynamics. The dynamics of EVD differ significantly across different geographic scales. Urbanized areas tend to experience outbreaks at a faster rate than rural areas because of the high population density, varied patterns of movement, and better medical infrastructure [18]. Moreover, the phenomenon of superspreader events, where a small proportion of the infected population is responsible for a suspiciously high number of secondary infections, adds a considerable amount of variability and complexity that challenges resolution by purely deterministic approaches [19, 20]. These characteristics need modeling methods that combine large-scale spatial dynamics with fine-scale spatio-temporal transmission patterns [21]. Although traditional modeling techniques such as the SI, SIR, and SEIR frameworks form the basis of understanding, they all fundamentally relate to the Markovian assumptions which do not take into account memory and long-range chronological dependencies which are characteristic of EVD spread [22, 23]. These limitations are particularly clear when attempting to model the distributed incubation periods, the decline of vaccine-induced immunity, and the heterogeneous contact patterns during the real-world epidemic spread [11, 24]. The SEIR model used by [25] and [26], and the SIHDP model employed by [27] and in [28] SEIHFDBR are notable developments, however, they are still limited in the integer-order nature of their models. Fractional calculus has long shown the potential to model memory effects and other non-local connections, and thus provides a framework well-intentioned of study to overcome some of the noted deficiencies: [29]. Unlike integer-order derivatives, which always adopt Markovian processes, fractional operators capture the distributed suspensions which are characteristic of epidemiological systems, such as extended incubation period, waning immunity, and environmental reservoirs of pathogens [30]. Evidence of progress in the modeling of infectious diseases for which data are available and for which fractional calculus applies exhibits greater concord with experimental indication, especially for diseases with long-range chronological correlations and power-law decline in the efficiency of involvements [31, 32]. The non-Markovian nature of some fractional derivatives offers a better-suited mathematical framework for modeling the transmission of disease in and among structured populations [33]. In addressing challenges, some recent studies have included various extensions. For example, [34] included pathways for environmental contamination in their SIRPD model with respect to the Ebola pathogens in the environment. In [35] worked on the decomposition of the deceased compartment by distinguishing dead but not buried, and dead who are buried. Control optimization has been studied by, for instance, [36] on optimal combinations of vaccination and treatment, and [37] on educational campaigns in a two-patch model. Media awareness campaigns have been studied by [38] and [39] and have shown the importance of information broadcasting in controlling a disease. The link of Machine Learning algorithms with Process-based model and models of machine learning with traditional epidemiological forecasting is a new concept as noted by [40]. Predictable models afford a theoretical understanding of epidemiology through biological lenses and as such are weak at capturing sophisticated structures non-linear in nature that are defined with basic compartmental approaches [41]. Recent activities in hybrid frameworks combining the Physics-based model interpretability and the neural networks’ model complexity to capture epidemiological dynamics have demonstrated improved predictive performance [42]. But, as advanced computational techniques apply to EVD, there are gaps such as real time parameter estimation from EVD noise where EVD observation data is incomplete [43]. The development of neural networks with physics-informed neural network (PINN) and epidemiology constraints could bridge this gap [44]. However, modern modeling system contexts still have three main shortcomings. One, they have a limited representation of the memory effects of fractional-order derivatives, which is crucial for modeling the distributed intervals in the progression of a disease and the corresponding delay of the immune response. Two, there is still a lack of fully integrating optimal control theory combined with advanced multi-compartment models that summarize the entirety of the EVD transmission dynamics. Three, here is still insufficient use of machine learning for the estimation of parameters for complicated epidemiological models which reduces the flexibility and accuracy of the models in current practice. The application of optimal control theory is underdeveloped yet central to the design of effective intervention and control strategies for the economic epidemiology systems. Identifying optimal control strategies is a challenging mathematical problem. It requires temporally-resolved control strategies and sophisticated numerical methods for high dimensional, nonlinear systems with memory [45]. Also, minimizing the case count, death, and economic loss for epidemic reduction response requires Pareto-optimal strategies that balance the contradictory objectives [46]. The recent, rapid advancements in adaptable computational optimal control technologies—particularly those dealing with the direct collocation of the control function and the system trajectory and the Hamiltonian-based methods are methods of great attention in computational biologically structured control [47]. The development of fractional calculus has presented new ways to tackle these problems [29]. The Caputo fractional derivative approaches non-local chronological dependencies and memory effects with its non-Markovian structure, which mathematically captures epidemiological data [48]. Preliminary work connecting fractional-order models focused on EVD, such as [49] with the Atangana-Baleanu-Caputo operators and [50] with the nonlinear time-fractional mathematical model. Despite these advances, there is still much to be anticipated in the overlap among these fields, mainly control theory, machine learning, and the other works on the theme. This research discourses these important gaps by revolutionary a new computational framework that combines fractional-order dynamics, multi-pathway transmission architectures, and control optimization, which is enhanced by machine learning. Building upon the preceding work by [34] which did not include exposed compartments, hospitalizations, and safe burial practices, I also consider reinfection through the immunological studies by [16] and [17]. This research directs numerous contributions and novelties:
@@
  • Comprehensive mathematical analysis signifying the well-posedness of the system through Banach’s fixed-point theorem, the basic reproduction number of the model $\mathcal{R}_{0}$ exogenously resulting through the next generation matrix, and the proof of global stability of the system through fractional Lyapunov methods [48].
  • Derivation and providing an analytical solution for a multi-objective optimal control problem with four intervention strategies (personal-protection, vaccination, treatment, and safe burial) via Pontryagin’s Maximum Principle.
  • Establishing a Disease-Informed Neural Network (DINN) framework for simultaneous solution discovery and parameter estimation which inserts physical constraints for better predictive performances.
  • Detailed cost-effectiveness analysis for optimal intervention portfolios for public health choices.
This model potentially assistances public health authorities in responding to outbreaks more effectively by evidence-based intervention, which can be game-changing in undertaking and enhancing the efficiency of outbreak responses. Integrating mathematical modeling with computational intelligence, this study attempts to productive infectious disease interventions through addressing the organized response and gaps in the current outbreak response system. This paper proceeds as follows. In Section 2, I discuss the elementary ideas of fractional calculus which is key to developing the model. In Section 3, I develop an eight-compartment, fractional-order Ebola transmission model. Section 4 contains the in-depth mathematical theory regarding the model’s well-posedness, equilibrium, and stability over the system. Section 5 conducts global sensitivity analysis of the dominant parameters through Latin Hypercube Sampling and Partial Rank Correlation Coefficients. In Section 6, I formulate and solve the optimal control problem based on Pontryagin’s Maximum Principle. Section 7 incorporates the neural network model concerning disease and the results obtained in the corresponding numerical analysis. In the closing Section 8, I analyze the outcome of the model along with its drawbacks and suggest some directions for further research.
@@ 2 Preliminaries
Definition 2.1 (Caputo Fractional Derivative)
Let $\alpha\in(0,1]$ and $u\in C^{1}([0,T])$. For any function $u$ which is continuously differentiable in the closure $[0,T]$ the Caputo fractional derivative of order $\alpha$ is defined in [29] as follows:
@@
Definition 2.2 (Mittag-Leffler Function)
The two-parameter Mittag-Leffler function is defined by the series [29]:
$E_{\alpha,\beta}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(\alpha k+\beta)},\quad\alpha,\beta>0,\,z\in\mathbb{C}.$
This function generalizes the exponential function and is crucial for constructing solutions to fractional differential equations, playing a role analogous to the exponential in integer-order systems.
Definition 2.3 (Fractional Lyapunov Stability)
@@
$\|y(t)-y^{*}\|\leq\left[m(y(0)-y^{*})E_{\alpha,1}(-\lambda t^{\alpha})\right]^{b},$
@@
Definition 2.4 (Physics-Informed Neural Network Loss)
For a system governed by $^{C}D_{t}^{\alpha}\mathbf{Y}(t)-\mathbf{F}(t,\mathbf{Y}(t))=0$, the loss function for a Physics-Informed Neural Network (PINN) $NN(t;\theta)$ approximating $\mathbf{Y}(t)$ is defined as:
$\mathcal{L}(\theta)=\mathcal{L}_{data}+\mathcal{L}_{physics},$
where $\mathcal{L}_{data}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}|NN(t_{i};\theta)-\mathbf{Y}_{i}|^{2}$ enforces data fidelity, and $\mathcal{L}_{physics}=\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}|{}^{C}D_{t}^{\alpha}NN(t_{j};\theta)-\mathbf{F}(t_{j},NN(t_{j};\theta))|^{2}$ enforces the physical law described by the fractional differential equation.
Theorem 2.1 (Stability Condition for Fractional Systems)
Let $\alpha\in(0,1]$. An equilibrium point $y^{*}$ of the autonomous system $^{C}D_{t}^{\alpha}y(t)=f(y(t))$ is asymptotically stable if all eigenvalues $\lambda_{i}$ of the Jacobian matrix $J=\partial f/\partial y|_{y^{*}}$ satisfy the condition [48]:
@@
3 Fractional-Order Ebola Transmission Dynamics
We formulate a Caputo fractional-order model ($\alpha\in(0,1]$) which incorporates the multi-pathway transmission, memory effects, and time-varying interventions of Ebola. The governing system is given as follows.
@@ 3 Fractional-Order Ebola Transmission Dynamics
The force of infection is given by:
@@
and for notational simplicity, we define:
@@
The model parameters include: natural mortality ($\mu$) and recruitment ($\Lambda$); vaccination rate $v(t)$ with waning immunity ($\omega$); and force of infection $\lambda(t)$ driving exposures. Exposed individuals ($E$) progress to infectious states at rate $\sigma$, splitting into symptomatic ($I_{s}$, proportion $p$) and asymptomatic ($I_{a}$, 1p1-p) pathways. Symptomatic cases face hospitalization ($h_{s}$), disease-induced mortality ($\delta_{s}$), or recovery ($\gamma_{s}$). Asymptomatic and hospitalized individuals recover at rates $\gamma_{a}$ and $\gamma_{h}$ respectively, with hospital mortality ($\delta_{h}$). Deceased individuals ($D$) transmit infection until safe burial ($\mu_{d}$). Transmission coefficients $\beta$, $\eta_{a}$, and $\eta_{d}$ govern infectivity from symptomatic, asymptomatic, and post-mortem sources. The system is governed by:
@@
In this Caputo fractional formulation, the system in (2) captures the dynamics of Ebola transmission in an eight-dimensional frame. This construction involves ‘epidemic’ components such as: Susceptible ($S$), Vaccinated ($V$), Exposed ($E$), Symptomatically ($I_{s}$) and Asymptomatically ($I_{a}$) infectious, Hospitalized ($H$), deceased ($D$), and Recovery ($R$) stages, thus integrating critical structure of underlying dynamics. Furthermore, the model participates the fractional-order system of differential equations, thus inserting memory effects.
@@
Table 1: Epidemiological state variables and parameter estimates for the fractional-order Ebola model.
Variable Epidemiological Definition Estimated Value
$S(t)$ Susceptible individuals
$V(t)$ Vaccinated individuals
$E(t)$ Exposed (infected, not infectious)
$I_{s}(t)$ Symptomatic infectious individuals
$I_{a}(t)$ Asymptomatic infectious individuals
$H(t)$ Hospitalized/isolated individuals
$D(t)$ Deceased (infectious pre-burial)
$R(t)$ Recovered individuals
$N(t)$ Total living population
$\beta$ Transmission rate 0.287
$\eta_{a}$ Relative infectiousness (asymptomatic) 0.523
$\eta_{d}$ Relative infectiousness (deceased) 0.734
$\sigma$ Incubation rate (day-1) 0.094
$p$ Proportion symptomatic 0.712
$\gamma_{s}$ Recovery rate (symptomatic) 0.068
$\delta_{s}$ Death rate (symptomatic) 0.103
$h_{s}$ Hospitalization rate 0.312
$\gamma_{a}$ Recovery rate (asymptomatic) 0.089
$\delta_{h}$ Death rate (hospitalized) 0.067
Table 2: Parameter ranges and biological interpretation for the fractional-order Ebola model.
Symbol Biological Interpretation Range Source
$\Lambda$ Recruitment rate 100-1000 [1]
$\mu$ Natural mortality rate $3.5\times 10^{-5}$ [51]
$c(t)$ Contact rate function 0.4-1.2 [9]
$v(t)$ Vaccination rate 0.005-0.08 [52]
$\varepsilon$ Vaccine efficacy 0.85-0.95 [52]
$\omega$ Waning immunity rate 1/365-1/270 [53]
$\mu_{d}$ Burial rate 0.15-0.35 [54]
$\alpha$ Fractional order 0.75-0.95 [29]
Figure 2: Model validation: Curve fitting to confirmed Ebola cases ($R^{2}=0.96$, RMSE = 23.45).

### v3

@@ Abstract
Ebola virus disease is a severe, often fatal infectious disease transmitted through direct contact with infected bodily fluids and contaminated surfaces. This study addresses the serious need for integrated predictive and control frameworks by developing a unified approach that connects intelligent outbreak prediction systems with mathematical models of Ebola transmission. We propose a novel nonlinear fractional-order eight-compartment SVEIsIaHDR model that integrates control measures, personal protection, vaccination, treatment, and safe burial to analyze their effectiveness in reducing Ebola transmission and improving outbreak prediction. Mathematical validation establishes the model’s boundedness, non-negativity, and well-posedness. We derive the basic reproduction number R0 via the next-generation matrix, identify disease-free and endemic equilibrium, and conduct stability analyses. Sensitivity analysis reveals R0’s primary sensitivity to transmission rate $\beta$ (+1.00), incubation period $\sigma$ (+0.6250), and infectivity of the dead $\eta_{d}$ (+0.4678). An optimal control framework minimizes disease management costs (4 dollars per infection). Numerical simulation by the RK45 method demonstrates combined treatment and safe burial controls 86.5% of target morbidity and mortality, with safe burial being the most effective single measure. We develop a disease-informed neural network (DINN) achieving 99.87% prediction accuracy ($R^{2}=0.9987$) with minimal error (NMSE $10^{-4}$ to $10^{-5}$). The DINN accurately recovers 13 of 17 transmission parameters with $9\%$ error, establishing a new interpretability benchmark. This integrated framework provides public health officials with a practical tool to test interventions and predict outbreaks through model-data integration.
Keywords: Ebola virus disease, fractional-order compartmental model, basic reproduction number ($\mathcal{R}_{0}$), optimal control, sensitivity analysis, stability analysis, outbreak prediction, disease-informed neural network (DINN).
@@ 1 Introduction
A serious infectious disease, Ebola Virus Disease (EVD) is transmitted to people by reservoirs of animals such as fruit bats. When the Ebola virus transmits from animal reservoirs into human populations, an outbreak occurs [1, 2]. The illness has historically significant case fatality rates, with outbreaks ranging from 25% to 90% [3]. It was first reported in the Democratic Republic of the Congo in 1976; the name Ebola was given due to the Ebola River [4]. With 28,646 confirmed infections and 11,323 deaths recorded, the most catastrophic epidemic was in West Africa (2014–2016), with almost a 50% death ratio [5]. Fever, weakness, diarrhea, and, in extreme situations, hemorrhagic signs are common in infected people. There is a major issue in that these viruses’ incubation periods differ, sometimes 2 days and sometimes 21 days; infected individuals cannot spread the virus throughout the incubation phase [6]. Fruit bats are considered to be the main natural source of infection in humans, and non-human primates serve as end hosts, maintaining cycles of transmission that can result in major outbreaks [7, 8]. Controlling measures are complicated, particularly frequent contact paths that boost the spread of EVD. Being in contact with infectious body fluids, such as blood or excretions, results in subsequent human-to-human transmission. Contact with deceased individuals during traditional burial practices is a crucial and widely recognized route of transmission that has been a significant facilitator of prior epidemics [9, 10]. In addition, the virus might continue to spread indirectly by remaining on contaminated surfaces. Due to the absence of apparent symptoms, various demographic categories, in particular people who participate in burial customs and asymptomatic individuals, have a possibility to spread infection throughout communities. Therefore, it is difficult for standard mathematical models to capture the EVD dynamics of transmission, as they have to immediately identify these various and context-specific threats [11, 12]. Ebola transmission is rendered much more difficult by the potential for reinfection. After recovery, immunity may not be total or long-lasting because protective antibodies may gradually decline in survivors [13, 14]. Individuals who have previously been infected can test positive for the virus years later, indicating potential re-exposure and infection, according to data from previous outbreaks [15, 16]. Additionally, research on populations that have been exposed frequently has discovered people who display no symptoms but do not have detectable antibodies, underscoring the heterogeneity in immune response [17, 18]. Now, we focus on a few equivalent mathematical models of EVD that have already been proposed by other researchers. Fractional calculus has been applied to infectious disease modeling in a number of research investigations, displaying more consistency with observed outbreak data than traditional approaches [19, 20]. Fractional calculus has long shown the potential to model memory effects and other non-local connections and thus provides a well-intentioned framework of study to overcome some of the noted deficiencies [21]. Unlike integer-order derivatives, which always adopt Markovian processes, fractional operators capture the distributed suspensions that are characteristic of epidemiological systems, such as extended incubation periods, waning immunity, and environmental reservoirs of pathogens [22]. Evidence of progress in the modeling of infectious diseases for which data are available and for which fractional calculus applies exhibits greater concord with experimental indication, especially for diseases with long-range chronological correlations and power law decline in the efficiency of involvements [23, 24]. The non-Markovian nature of some fractional derivatives offers a better-suited mathematical framework for modeling the transmission of disease in and among structured populations [25].The SI model, such as [26, 27], the SEIR model [28, 29], the SEIRD model [30], and the SEIRHD model [31, 32, 33, 34], where $S$, $E$, II, $H$, $R$, and $D$ represent susceptible, exposed, infectious, hospitalized, recovered, and dead/deceased compartments, respectively [35], separated the Dead/Deceased ($D$) compartment into two compartments, buried deceased ($R_{D}$) and unburied dead ($F$). Similarly, in their SIRPD model, where $P$ represents environmental contamination, [26] included EVD transmission through contaminated surroundings of deceased patients. While models like the SEIHFDBR in [36] represent notable developments, they remain limited by their integer-order formulation. Understanding how to effectively control the spread of infection and facilitate decision-making during an epidemic is one of the applications of mathematical modeling. For example, [37] presented an SIR model-based optimum control issue with treatment and immunization as possible control strategies. They looked into the most effective vaccination and therapy combinations to lower the cost of those preventive measures. In [38] investigated various diseases transmitted by insects that directly infect the host population, such as dengue fever, malaria, and the West Nile virus, using optimal control analysis. In order to examine the transmission of Ebola, [39] designed a model. They evaluated the effects of various preventative measures such as vaccination, public awareness campaigns, early identification campaigns, improved hospital sanitation, separation of people with infections, and geographical constraints. Another field of research is the integration of machine learning with traditional epidemiological models. When handling noisy or incomplete surveillance information, physics-informed neural networks (PINNs) demonstrate particular potential for parameter estimation and prediction [40, 41, 42]. A comprehensive analysis of the last five years’ worth of literature, however, reveals enduring shortcomings and disparate study directions. Although fractional models have been created, they are frequently not integrated with intricate multi-pathway transmission designs specific to EVD, for instance. Likewise, machine learning applications often function independently of mathematical models, which limits their understanding and generalization, and optimal control formulations usually use simplified models that do not fully capture the complexity of transmission dynamics [43]. Nevertheless, despite these developments, there are still important research gaps in present literature. Firstly, there is a gap between the full representation of all recognized EVD transmission channels (symptomatic, asymptomatic, and burial customs) and fractional-order models, which account for memory effects. There is a methodological gap given that detailed transmission models generally employ integer-order approaches, whereas the majority of fractional models simplify transmission dynamics. Second, the application of optimal control theory to high-dimensional fractional systems that reflect EVD dynamics has been insufficient. One major challenge that has limited practical applicability is the mathematical complexity of determining optimal control policies for such systems, especially when including numerous intervention strategies with different costs and effectiveness. [44, 45]. Third, the majority of applications currently only combine biological process-based models with data-driven machine learning techniques. Hybrid frameworks have been proposed, but they typically lack the seamless combination needed for real-time parameter estimation and prediction during current outbreaks [46]. This study develops a novel computational framework with the particular objectives that follow in order to address these connected challenges. The practical objective is to provide public health decision-makers an efficient computational tool that can forecast disease paths with measured uncertainty, optimize allocation of resources for maximum impact, and simulate outbreak scenarios under different intervention strategies. The following is a list of this study’s primary innovations and contributions:
@@
  • We established the optimal control model that would help to reduce the cost of managing the disease by considering four intervention measures that included personal protection, vaccination, treatment and safe burial. Our theoretical analysis by applying Pontryagin’s Maximum Principle and numerical simulation using the RK45 method permitted identifying the cost-effective strategies, and the level of their effectiveness in terms of transmission and mortality reduction.
  • Motivated from physics-informed neural networks, we designed a new framework for disease-informed neural networks (DINN). This computational approach achieves high prediction accuracy across all disease stages and demonstrates strong parameter recovery capability, providing an interpretable tool for outbreak forecasting and intervention assessment.
This work significantly improves both academic research and healthcare practice by successfully addressing the existing gap between theoretical modeling and actual application. It offers investigators a new methodological framework that shows how machine learning, optimal control theory, and fractional calculus may be successfully used for the scientific study of infectious diseases. This paper proceeds as follows. In Section 2, I discuss the elementary ideas of fractional calculus, which is key to developing the model. In Section 3, I develop an eight-compartment, fractional-order Ebola transmission model. Section 4 presents the complete mathematical analysis of the model, including proofs of existence and uniqueness of solutions, identification of the invariant region and attractivity, derivation of the basic reproduction number ($\mathcal{R}_{0}$), and both local and global stability analyses for the disease-free and endemic equilibrium points. Section 5 performs a global sensitivity analysis that attempts to quantity and prioritize the impact of key model parameters on the disease transmission dynamics. Section 6 establishes and solves optimal control problem through the Maximum Principle of Pontryagin founding the most effective intervention strategies. In section 7, the formation of a disease-informed neural network (DINN) model along with the relevant numerical data is developed and fully analyzed. Lastly, in Section 8 the main findings are synthesized, the methodological limitations are critically discussed, and future research directions outlined.
@@ 2 Preliminaries
Definition 2.1
Let $\alpha\in(0,1]$ and $u\in C^{1}([0,T])$. For any function $u$ which is continuously differentiable in the closure $[0,T]$ the Caputo fractional derivative of order $\alpha$ is defined in [21] as follows.
@@
Definition 2.2
@@
$\|y(t)-y^{*}\|\leq\left[m(y(0)-y^{*})E_{\alpha,1}(-\lambda t^{\alpha})\right]^{b}.$
@@
Lemma 2.1
Let $\alpha\in(0,1]$. An equilibrium point $y^{*}$ of the autonomous system $^{C}D_{t}^{\alpha}y(t)=f(y(t))$ is asymptotically stable if all eigenvalues $\lambda_{i}$ of the Jacobian matrix $J=\partial f/\partial y|_{y^{*}}$ satisfy the condition [47].
@@
3 Model formulation
We formulate a Caputo fractional-order model ($\alpha\in(0,1]$) that incorporates the multi-pathway transmission, memory effects, and time-varying interventions of Ebola. The governing system is given as follows.
@@ 3 Model formulation
The force of infection is given by
@@
and for notational simplicity, we define the model as
@@
The model parameters include natural mortality ($\mu$) and recruitment ($\Lambda$), vaccination rate $v(t)$ with waning immunity ($\omega$), and force of infection $\lambda(t)$ driving exposures. Exposed individuals ($E$) progress to infectious states at rate $\sigma$, splitting into symptomatic ($I_{s}$, proportion $p$) and asymptomatic ($I_{a}$, 1p1-p) pathways. Symptomatic cases face hospitalization ($h_{s}$), disease-induced mortality ($\delta_{s}$), or recovery ($\gamma_{s}$). Asymptomatic and hospitalized individuals recover at rates $\gamma_{a}$ and $\gamma_{h}$, respectively, with hospital mortality ($\delta_{h}$). Deceased individuals ($D$) transmit infection until safe burial ($\mu_{d}$). Transmission coefficients $\beta$, $\eta_{a}$, and $\eta_{d}$ govern infectivity from symptomatic, asymptomatic, and post-mortem sources.The system is governed by,
@@
In this Caputo fractional formulation, the system in (2) captures the dynamics of Ebola transmission in an eight-dimensional frame. This construction involves ‘epidemic’ components such as susceptible ($S$), vaccinated ($V$), exposed ($E$), symptomatically ($I_{s}$) and asymptomatically ($I_{a}$) infectious, hospitalized ($H$), deceased ($D$), and recovery ($R$) stages, thus integrating the critical structure of underlying dynamics. Furthermore, the model participates in the fractional-order system of differential equations, thus inserting memory effects.
Table 1: Parameters for the fractional-order Ebola model.
Symbol Description Value Range Source
$\beta$ Transmission rate 0.287 0.2–0.4 [48]
$\eta_{d}$ Deceased infectiousness 0.734 0.5–0.9 [48]
$\eta_{a}$ Asymptomatic infectiousness 0.523 0.3–0.7 Assumed (Fitted)
$\sigma$ Incubation rate (day-1) 0.094 0.05–0.15 [49]
$p$ Symptomatic proportion 0.712 0.5–0.8 [49]
$\gamma_{s}$ Symptomatic recovery (day-1) 0.068 0.05–0.1 [26]
$\gamma_{a}$ Asymptomatic recovery (day-1) 0.089 0.07–0.12 [26]
$\delta_{s}$ Symptomatic death (day-1) 0.103 0.05–0.15 [50]
$\delta_{h}$ Hospitalized death (day-1) 0.067 0.03–0.09 [50]
$h_{s}$ Hospitalization rate 0.312 0.2–0.4 [51]
$\Lambda$ Recruitment rate 100 100–1000 [52]
$\mu$ Natural mortality (day-1) $3.5\times 10^{-5}$ [52]
$v(t)$ Vaccination rate 0.005–0.08 Assumed (Control)
$\varepsilon$ Vaccine efficacy 0.85–0.95 Assumed (Control)
$\omega$ Waning immunity (day-1) 0.0027–0.0037 Assumed
$\alpha$ Fractional order 0.85 0.75–0.95 Assumed (Model)